OpenStudy (anonymous):

What would be the discriminant of h(x)=x^2-2x-35? How many complex solutions will it have? And can it be solved by the Quadratic Formula. Yes, this is easy stuff but its been so long that i forgot how to do it. Thanks

OpenStudy (anonymous):

Your function h has two real roots, since \[ h(x)=x^2-2 x-35=(x-7) (x+5) \]

OpenStudy (anonymous):

If you want to do it using the quadratic formula, \[ \Delta = b^2 - 4 a c= 4+140=144\\ \sqrt {\Delta}=12\\ x=\frac { 2\pm 12 }{2}=1\pm 6\\ x=7\\ x=-5 \]